OpenStudy (anonymous):

3u/u^2-1 + u/u-1, simplify

OpenStudy (anonymous):

3u/(u-1)(u+1) + u/u-1=3u/(u-1)(u+1) +u(u+1)/(u-1)(u+1)=3u+u^2+u/u^2-1=u^2+4u/u^2-1